OpenStudy (anonymous):

Simplify the expression. q^6/q^10 1/q^4 q^16 q^4 1/q^16

OpenStudy (anonymous):

\[\frac{ x^m }{ x^n } \implies x^{m-n}\] try it out yourself now and see what you get :)

OpenStudy (anonymous):

saying 1/4 q

OpenStudy (anonymous):

\[\frac{ 1 }{ q^4 }\] if you mean this, that's right

OpenStudy (anonymous):

\[\frac{ q^6 }{ q^{10}} \implies \frac{ 1 }{ q^{10-6} } \implies \frac{ 1 }{ q^4 }\]

OpenStudy (anonymous):

Yaythankyou im getting it right then using this calculator thankyou kindly :)